# Estimating the Lensing Probability for Binary Black Hole Mergers in AGN disk by Using Mismatch Threshold

### v1

@@ Abstract
Stellar-mass binary black holes (BBH) may form, evolve, and merge within the dense environments of active galactic nuclei (AGN) disks, thereby contributing to the BBH population detected by gravitational wave (GW) observatories. Mergers occurring in AGN disks may be gravitationally lensed by the supermassive black hole at the AGN centre. The probability of such lensing events has been approximately estimated by using the Einstein criterion in previous work. However, a more reasonable approach to calculating the lensing probability should be based on whether the detector can distinguish the lensed GW waveform from the unlensed one. In this work, we calculate the lensing probability of LIGO sources embedded in AGN disk by relating threshold mismatch to the signal-to-noise ratio of the observed events. For the sensitivity of LIGO-Virgo-KAGRA O3 observation runs, our results indicate that the lensing probability is several times higher than previous estimates. If AGNs are indeed the primary formation channel for BBHs, we could quantify the probability of detecting the lensed GW events in such a scenario. The non-detections, on the other hand, will place stricter constraints on the fraction of AGN disk BBHs and even the birthplaces of BBH mergers.
]Dated: May 11, 2025
@@ I Introduction
In 2015, LIGO detected the first binary black hole (BBH) merger event, known as GW150914 [1, 2, 3]. Two years later, another significant event was observed – the merger of a binary neutron star, GW170817 [4]. This event was accompanied by the electromagnetic counterpart of GRB170817A [5, 6] and Kilonova AT 2017gfo/DLT17ck [7]. Gravitational waves (GWs) from two neutron star–black hole coalescences were also observed in January 2020 [8]. Up to date, about 90 confident GW detections have been made by LIGO-Virgo-KAGRA (LVK) [9, 10, 11, 12, 13]. As the number of detected GW events continues to rise, the formation channels of compact binary systems and the physical mechanisms behind these mergers have extensively studied.
BBHs can form, evolve and merge in the densest environments of AGNs [14, 15, 16, 17, 18]. Stellar-mass black holes in AGNs will become gravitationally bound to one another due to the effects of gaseous dynamical friction from the disk [19, 20, 21, 22, 23, 24, 25, 26, 27, 28] or stellar dynamical friction from the nuclear star cluster [29]. Stars [30, 31], star clusters [32, 33] and BBHs that are orbiting around the central SMBH may interact with the AGN disk and captured by it, thereby forming potential populations of GW or tidal disruption events [34, 35, 36]. These objects embedded within the disk may sink to the so-called migration trap region [37, 38, 39, 40] and form BBHs. Some of them may form a hierarchical triple system with the central supermassive black hole (SMBH) and merge in the manner of Kozai–Lidov mechanism [41, 42].
Previous work obtained a significant merger rate of BBHs on the AGN disk through Monte Carlo simulations [43]. A natural consideration is that the GW from BBH mergers can be lensed by the SMBH when the BBH is located behind the SMBH. Inspecting the lensing probability and comparing with the detection results of lensed GW events could put strong constraint on the number of such systems and the merger sites of BBH in the disk.
For such a BBH+SMBH lensing system, the criterion is often approximately characterized by the Einstein radius $\theta_{\rm Ein}$ in previous work [44, 45]. [45] find that, with $\sim 100$ observations, the lensing probability of observing at least one GW 190521-like is about $14\%$. However, a more reasonable approach to calculating the lensing probability should be based on whether the detector can distinguish the lensed GW waveform from the unlensed one [46, 47, 48, 49, 50]. The criterion should depend on the sensitivity of the detector and the SNR of the detected GW events. In this work, we calculate the lensing probability of LIGO sources embedded in AGN disk by relating threshold mismatch to the SNR.
This paper is organized as follows. In Section II, we present our model. In Section III, we present our results on the mismatch arising from gravitational lensing and the lensing probability. Conclusions and discussions are given in Section IV.
@@ II Gravitational Lensing of BBH mergers in AGN disk and mismatches
Figure 1: Schematic diagram of the lensing system. Figure 1 $(a)$ showing a BBH (Source $S$) orbiting an SMBH (Lens object $L$). The gray-shaded region indicating the orbital plane of the source. When the BBHs are located in the conical region on the right, they will form a lensing system with the central SMBH and the observer $O$ on the left. $D_{\rm L}$, $D_{\rm S}$ and $D_{\rm LS}$ represent the distance from the observer to lens, the observer to source and the lens to source, respectively. $\beta$ is the angular position of the source. The position of the source in the disk is represented by $\phi$, $\iota$, and $a$, where $\phi$ denotes the orbital phase, $\iota$ indicates the angle between the source’s orbital angular momentum ($\hat{L}_{\rm S}$) and optical axis ($x$-axis), while $a$ represents the semi-major axis of the source’s orbit. Several cones with opening angles of $\phi_{\rm th,max}$ are also plotted to depict the scenario when lensing of GW will likely occur. We get Figure 1$(b)$ by setting $\iota=90^{\circ}$. $\phi_{\rm th,max}\approx 6^{\circ}$ and $10^{\circ}$ corresponds to the maximum threshold orbital phase when the SNR are $8$ and $15$. The blue solid angle of $\phi_{\rm Ein}\approx 3^{\circ}$ is given by Einstein criterion.
@@
Figure 2: Contour plot of the characteristic frequency, $f_{M_{\rm{Lz}}}$. Wave-optics dominated zone: bottom left region, where $f_{\rm{GW}}\sim f_{M_{\rm{Lz}}}$. The frequency $f_{M_{\rm{Lz}}}=10$ $\rm{Hz}$ corresponds to the lower limit of the LIGO-Virgo sensitivity band. Geometrical-optics zone: upper right region, where $f_{\rm{GW}}\gg f_{M_{\rm{Lz}}}$. For reference, we have plotted the Einstein criterion determined phase angles $\phi_{\rm{Ein,1000}}$ in dashed black line, when $\iota=90^{\circ}$. As seen from the dashed black line in the figure, the Einstein criterion related orbital phase fall entirely within the geometric optics limit.
@@
The geometric configuration of the gravitational lensing effect of the GW adopted in this work is depicted in Figure 1 ($a$). Here, the source is the BBH, which is orbiting the SMBH that is considered as the lensing object $L$. The reference frame $(x,y,z)$ considered here is centered on the lensing object $L$. The $x$-$\rm{axis}$ is defined by the line of sight between the observer and the lensing object $L$. The $x$-$y$ plane is defined by the plane containing both the $x$-$\rm{axis}$ and the source’s orbital angular momentum vector $\hat{L}_{\rm S}$ around the object $L$ [51, 47, 52], the angle between the $y$-$\rm{axis}$ and $\hat{L}_{\rm S}$ is $90^{\circ}-\iota$. The source location can thus be written as
$\left\{\begin{array}[]{lr}x(a,\phi,\iota)=a\cos\phi\sin\iota,&\\ y(a,\phi,\iota)=a\cos\phi\cos\iota,&\\ z(a,\phi)=-a\sin\phi,&\end{array}\right.$ (1)
@@
where $a$ is the semi-major axis of the source, $\phi$ is the orbital phase of the source and $\iota$ is the inclination angle between the line of sight and the orbital angular momentum of the source $\hat{L}_{\rm S}$, with $cos(\iota)=\boldsymbol{x}\cdot\boldsymbol{\hat{L}_{\rm S}}$, respectively. From Figure 1 ($a$), Equations (1) and (3), we can see that the angular source position $\beta$ is determined by $a,\phi,\iota$. Specifically, when $\iota=0^{\circ}$, this corresponds to a face-on disk. The angular position of the source is $\beta=a/D_{\rm L}$, independent of the source’s orbital phase $\phi$, and the lensing effect disappears in such a configuration. On the other hand, for a perfectly aligned lensing system, we will have $\iota=90^{\circ}$, $\phi=0^{\circ}$ and $\beta=0^{\circ}$. The dimensionless angular source location of point mass lens model [53] can be expressed as
@@
Following [54] and [52], the transmission factor of wave optics solution that depends on coordinate parameters have the form of
@@
×$\times\Gamma(1-i\frac{w}{2})_{1}F_{1}(i\frac{w}{2},1;i\frac{w\eta^{2}}{2}),$ (7)
where $\Gamma$ is complex gamma function, ${}_{1}F_{1}$ is confluent hypergeometric function of the first kind, and
@@
is the dimensionless frequency. $\phi_{m}(a,\phi,\iota)$ is defined as
@@
Following [55], we use the relation $f_{M_{\rm Lz}}\equiv 1/\Delta t$ to determine which region of the parameter space—Geometrical-optics regime, wave-dominated zone, or long-wavelength regime—our sources reside in. Following [48], we make a smooth connection from wave optical solution (see Equation (7)) to geometrical optic limits (i.e. Equation (11)).
For a given unlensed waveform template $h_{\rm UL}(f)$, the lensed waveform $h_{\rm L}(f)$ is defined as the product of the transmission factor and the unlensed waveform, as
$h_{\rm L}(f,a,\phi,\iota,M_{\rm Lz})=F(f,a,\phi,\iota,M_{\rm Lz})h_{\rm UL}(f).$ (12)
The match ($\mathcal{M}$) between these two waveforms provides a method to quantify their similarity, which has the value from zero to unity [56, 57]. The match between lensed and unlensed waveforms are
@@
[…]
is the noise weighted inner product between lensed and unlensed waveforms, with $S_{\rm n}(f)$ as the one-sided noise power spectral density. As a result, the difference between the lensed and unlensed waveforms is represented by
@@
The mismatch threshold $\epsilon_{th}$ is used to determine whether the two waveforms can be distinguished [58, 59, 56, 60, 49],
@@
where parameter $\phi_{\rm{th}}$ represents the threshold orbital phase corresponding to the mismatch threshold $\epsilon_{\rm th}$, and $\rho$ is SNR. For a lensing event with a typical detection of $\rho=10$, this means that $\epsilon\geq\epsilon_{\rm th}=0.005$ is an acceptable mismatch to indicate the difference between the lensed and unlensed. Therefore, we may use Equation (16) to estimate the lensing probability.
@@ III.1 Mismatch Arising from Lensing
In this section, we calculate the mismatch arising from the gravitational lensing effect, and estimate the lensing probability by using mismatch threshold $\epsilon_{\rm th}$.
[…]

### v2

@@ Abstract
Stellar-mass binary black holes (BBH) may form, evolve, and merge within the dense environments of active galactic nuclei (AGN) disks, thereby contributing to the BBH population detected by gravitational wave (GW) observatories. Mergers occurring in AGN disks may be gravitationally lensed by the central supermassive black hole. Such lensing presents a potential probe of BBH formation channels through its probability. In this work, we calculate the AGN lensing probability of LIGO sources embedded in AGN disk by relating mismatch threshold to the signal-to-noise ratio. The AGN lensing probability of GW190521-like event will be $\sim 3\%$ for O3 PSD, $\sim 6\%$ for Aplus PSD, and $\sim 33\%$ for ET PSD. If AGNs are indeed the primary formation channel for BBHs, we could quantify the probability of detecting the lensed GW events in such a scenario. The non-detections, on the other hand, will place stricter constraints on the fraction of AGN disk BBHs and even the birthplaces of BBH mergers.
]Dated:
@@ I Introduction
In 2015, LIGO detected the first binary black hole (BBH) merger event, known as GW150914 [7]. Two years later, another significant event was observed – the merger of a binary neutron star, GW170817 [8]. This event was accompanied by the electromagnetic counterpart of GRB 170817A [9, 50] and kilonova AT 2017gfo/DLT17ck [116]. Gravitational waves from two neutron star–black hole (BH) coalescences were also observed in January 2020 [12]. The network of ground-based gravitational-wave (GW) detectors, after four observing runs, has observed more than 300 events, most of which are merging BBHs [10, 11, 5, 14, 110, 70]. More GW events are expected to be detected in future observation runs and with new detectors such as LIGO-india [96], Deci-hertz Interferometer GW Observatory (DECIGO, [62]), Einstein Telescope (ET, [95]), and Cosmic Explorer (CE, [41]). This rapidly growing population motivates detailed studies of BBH formation channels.
The proposed scenarios include chemically homogeneous evolution of isolated binary stars [37], common-envelope ejection [90, 59], envelope expansion [108], chemically homogeneous evolution in a tidally distorted binary [37], evolution of triple or quadruple systems [18, 74], gravitational capture [87], dynamical evolution in open clusters [19] and dense star clusters [93], dynamical interaction in gas-rich nuclear regions [79], and the evolution of Population III binary stars [102], among others. However, the observation of BBH merger event GW231123 indicates that BBH masses are consistent with intermediate mass, and both black hole spins are consistent with very high spin parameters [1]. These two distinctive properties pose challenges to conventional formation channels.
Recent studies suggested that GW231123 may originate from active galactic nuclei (AGN channel). For example, Ref. [22] proposed that GW231123 could be formed via accretion onto Population III remnants in an AGN disk, whereas Ref. [38] suggested that it may arise from hierarchical mergers within AGN disks. Although electromagnetic counterpart observations could, in principle, provide constraints on the localization and environment of GW231123 [23, 30], whether such counterparts are detectable remains uncertain [78, 63, 122]. Thus, a single observational signature may be insufficient to uniquely identify the formation environment of BBHs, highlighting the need for multiple diagnostic approaches. In this context, we propose that AGN lensing may serve as a complementary diagnostic tool.
The basis for this proposal is that the complex environment of AGN may lead to diverse AGN lensing signatures, which in turn underpin their diagnostic potential. For example, BBHs prior to merger, while orbiting the central supermassive black hole (SMBH), may undergo repeated gravitational lensing events [34, 121], leading to periodic spikes in GW amplitude. As the BBH passes near caustics, its waveforms could be distorted due to diffraction [44]. The presence of a lensing object may also cause the GW to experience the gravitational spin Hall effect during propagation [119, 17, 88, 89], thereby altering its polarization [57]. Moreover, the stellar field [31, 103] surrounding the BBH and galaxy substructure [73, 104] may also play a role in shaping the AGN lensing signature.
Although the AGN lensing signatures are diverse, they are not beyond investigation. Previous work has established a framework for lensed GW signal search and identification [4, 2, 111], which provides a foundation for such investigations. Generally, each GW event is first identified by matched-filtering search pipelines [10, 11, 5, 14, 110, 70]. The assessment of potential lensing is performed subsequently [54, 35, 4, 60, 2, 111, 52, 28, 58], through the identification of lensing signatures [56, 42, 68] using model selection methods [61, 75, 55, 118]. Reference [81] employed a comprehensive approach combining Bayesian inference and matched-filtering analysis to systematically investigate lensing effects. Their results indicate that the maximum match across the $M_{\rm L}$$y$ plane effectively captures the similarity between lensed and unlensed waveforms. This motivates us to employ the match-filtering method to investigate and quantify the distribution of waveform deviations induced by AGN lensing across the AGN disk.
Existing research indicates that waveform deviations can be quantified using mismatch [71, 112]. With a reasonable construction of the AGN lensing geometry, it is possible to map the distribution of waveform deviations induced by AGN lensing across the disk via the mismatch. Unfortunately, even if such waveform deviations are quantified, they may not necessarily be measurable. Recent studies show that the mismatch can be related to an indistinguishability signal-to-noise ratio (SNR) [112]. Specifically, if the mismatch $\epsilon$ falls below a certain threshold (mismatch threshold hereafter), the intrinsic deviation between waveforms becomes smaller than the statistical uncertainty [112]. In such cases, the waveforms are considered indistinguishable [46, 47, 32, 48, 71, 72, 80]. This indistinguishability suggests two implications. First, AGN lensing effects need to be sufficiently pronounced to have a chance of being distinguished from other AGN-related effects. Second, for a merger event at a given disk location, if the AGN lensing-induced waveform deviation is indistinguishable, that event would not contribute to the overall AGN lensing probability. This offers a potential method for estimating the AGN lensing probability across the disk using the mismatch threshold and SNR.
GW lensing rates have been derived from optical depth, adopting the Einstein radius serving as a key criterion [33]. Using this approach, Ref. [66] found that, with $\sim 100$ observations, the lensing probability of observing at least one GW190521-like event is about $14\%$. However, the Einstein-radius criterion is not designed to reflect the specific distribution of BBHs within AGN disks, nor does it directly incorporate the effects of intrinsic binary parameters [69, 120]. Therefore, this motivates us to investigate the AGN lensing probability based on the mismatch threshold. This criterion, in turn, is shaped by several key factors: the source distance, the combination of binary mass, and the detector power spectral density (PSD), which constitutes the main subject of this study.
This work is organized as follows. In Section II, we present the geometric configuration of our AGN lensing model and match-filtering method. In Section III, we present calculations of the waveform mismatch induced by AGN lensing and its disk distribution, together with the estimation of AGN lensing probability. Conclusions and discussions are given in Section IV.
@@ II Gravitational Lensing of BBH mergers in AGN disk and mismatches
Figure 1: Schematic diagram of the lensing system. FIG. 1 $(a)$ showing a BBH (Source $S$) orbiting an SMBH (Lens object $L$). The gray-shaded region indicating the orbital plane of the source. When the BBHs are located in the conical region on the right, they will form a lensing system with the central SMBH and the observer $O$ on the left. $D_{\rm L}$, $D_{\rm S}$ and $D_{\rm LS}$ represent the distance from the observer to lens, the observer to source and the lens to source, respectively. $\beta$ is the angular position of the source. The position of the source in the disk is represented by $\phi$, $\iota$, and $a$, where $\phi$ denotes the orbital phase, $\iota$ indicates the angle between the source’s orbital angular momentum ($\hat{L}_{\rm S}$) and optical axis ($x$-axis), while $a$ represents the semi-major axis of the source’s orbit. Several cones with opening angles of $\phi_{\rm th,max}$ are also plotted to depict the threshold orbital phase. We get FIG. 1 $(b)$ by setting $\iota=90^{\circ}$. $\phi_{\rm th,max}\approx 6^{\circ}$ and $10^{\circ}$ corresponds to different semi-major axis. The blue solid angle of $\phi_{\rm Ein}\approx 3^{\circ}$ is given by Einstein criterion.
@@
Figure 2: Contour plot of the characteristic frequency, $f_{\rm{M}_{\rm{Lz}}}$. Wave-optics dominated regime: bottom left region, where $f_{\rm{GW}}\sim f_{\rm{M}_{\rm{Lz}}}$. The frequency $f_{\rm{M}_{\rm{Lz}}}=10$ $\rm{Hz}$ corresponds to the lower limit of the LIGO-Virgo sensitivity band. Geometrical-optics regime: upper right region, where $f_{\rm{GW}}\gg f_{\rm{M}_{\rm{Lz}}}$. For reference, we have plotted the Einstein criterion determined phase angles $\phi_{\rm{Ein,1000}}$ in dashed black line, when $\iota=90^{\circ}$. As seen from the dashed black line in the figure, the Einstein criterion related orbital phase fall entirely within the geometric optics limit.
@@
The geometric configuration of the AGN lensing effect of the GW adopted in this work is depicted in FIG. 1 ($a$). Here, the source is the BBH, which is orbiting the SMBH that is considered as the lensing object $L$. The reference frame $(x,y,z)$ considered here is centered on the lensing object $L$. The $x$-$\rm{axis}$ is defined by the line of sight between the observer and the lensing object $L$. The $x$-$y$ plane is defined by the plane containing both the $x$-$\rm{axis}$ and the source’s orbital angular momentum vector $\hat{L}_{\rm S}$ around the object $L$ [45], the angle between the $y$-$\rm{axis}$ and $\hat{L}_{\rm S}$ is $90^{\circ}-\iota$. The source location can thus be written as
[…]
@@
where $a$ is the semi-major axis of the source, $\phi$ is the orbital phase of the source and $\iota$ is the inclination angle between the line of sight and the orbital angular momentum of the source $\hat{L}_{\rm S}$, with $cos(\iota)=\boldsymbol{x}\cdot\boldsymbol{\hat{L}_{\rm S}}$, respectively. From FIG. 1 ($a$), Eq. (1) and (3), we can see that the angular source position $\beta$ is determined by $a,\phi,\iota$. Specifically, when $\iota=0^{\circ}$, this corresponds to a face-on disk. The angular position of the source is $\beta=a/D_{\rm L}$, independent of the source’s orbital phase $\phi$, and the AGN lensing effect may disappears in such a configuration. On the other hand, for a perfectly aligned AGN lensing system, we will have $\iota=90^{\circ}$, $\phi=0^{\circ}$ and $\beta=0^{\circ}$. The dimensionless angular source location of point mass lens model [101] can be expressed as
@@
Following Ref. [109], the amplification factor of the wave optic solution that depends on the coordinate parameters has the form of
@@
$\times\Gamma(1-i\frac{w}{2})_{1}F_{1}(i\frac{w}{2},1;i\frac{w\eta^{2}}{2}),$ (7)
where $\Gamma$ is the complex gamma function, $_{1}F_{1}$ is confluent hypergeometric function of the first kind, and
@@
is the dimensionless frequency. $\phi_{\rm m}(a,\phi,\iota)$ is defined as
@@
Following Ref. [81], we use the relation $f_{M_{\rm Lz}}\equiv 1/\Delta t$ to determine which region of the parameter space, the geometrical optical regime, the wave-dominated regime, or the long-wavelength regime, our sources reside in. Following Ref. [25], we make a smooth connection from wave optical solution (see Eq. (7)) to geometrical optic limits (i.e. Eq. (11)).
For a given unlensed waveform template $\tilde{h}_{\rm UL}(f)$, the lensed waveform $\tilde{h}_{\rm L}(f)$ is defined as the product of the amplification factor and the unlensed waveform, as
$\tilde{h}_{\rm L}(f,a,\phi,\iota,M_{\rm Lz})=F(f,a,\phi,\iota,M_{\rm Lz})\tilde{h}_{\rm UL}(f).$ (12)
The match ($\mathcal{M}$) between these two waveforms provides a method to quantify their similarity, which has a value of zero to unity [32]. The match between lensed and unlensed waveforms are
@@
$\langle h_{\rm L}|h_{\rm UL}\rangle\equiv 2\int\frac{\tilde{h}^{\ast}_{\rm L}(f)\tilde{h}_{\rm UL}(f)+\tilde{h}_{\rm L}(f)\tilde{h}^{\ast}_{\rm UL}(f)}{S_{\rm n}(f)}df\vskip 10.00002pt$ (14)
is the noise weighted inner product between lensed and unlensed waveforms, with $S_{\rm n}(f)$ as the one-sided noise power spectral density. As a result, the AGN lensing induced waveform deviation is represented by the mismatch between the lensed and unlensed waveforms
@@
The mismatch threshold $\epsilon_{th}$ is used to determine whether the two waveforms can be distinguished [46, 47, 32, 48, 71, 72, 80],
@@
where parameter $\phi_{\rm{th}}$ represents the threshold orbital phase corresponding to the mismatch threshold $\epsilon_{\rm th}$, and $\rho$ refers to the SNR. The square of the SNR of the lensed waveform is defined as
$\rho^{2}=4\int{\frac{|\tilde{h}_{\rm{L}}(f)|^{2}}{S_{\rm{n}}(f)}df}.$ (17)
@@ III Mismatch and Lensing Probability
In this section, given the dependence between the mismatch threshold and the SNR, we first calculate the SNR, then compute the mismatch arising from the AGN lensing effect, and finally estimate the lensing probability using the mismatch threshold $\epsilon_{\rm th}$.
III.1 Signal-to-Noise Ratio
In this subsection, we employ PyCBC [16, 15, 36, 114, 85, 84] to calculate the SNR. Our calculations primarily consider the following dependence: source distance, different mass combinations, and three different detector sensitivities.
The mass of sources observed by LIGO range from several solar masses to few hundred solar masses, with the merger rate following a broken power-law distribution [3]. Accordingly, we select $20M_{\odot}+20M_{\odot}$ and $35M_{\odot}+30M_{\odot}$ as representative examples for our calculations. Furthermore, considering that higher mass events may result from hierarchical mergers [49], we also adopt the mass combination corresponding to the GW190521-like event (e.g $85M_{\odot}+66M_{\odot}$) [3]. Regarding the PSDs, our focus lies on how improvements in PSD affect the SNR, as well as the potential enhancement in SNR from future multi-detector joint observations. Therefore, we use the aLIGOaLIGOO3LowT1800545 PSD [13] (O3 PSD hereafter) as a baseline reference and compute the variations in SNR under both the aLIGOAPlusDesignSensitivityT1800042 PSD (Aplus PSD hereafter) [21] and the EinsteinTelescopeP1600143 PSD (ET PSD hereafter) [40] scenarios.
Figure 3: SNR as a function of source distance. The purple, orange and green lines represent O3 PSD, Aplus PSD, and ET PSD, respectively. The solid, dashed and dot-dashed lines correspond to three mass combinations, i.e., $m_{1}=m_{2}=20M_{\rm{\odot}}$, $m_{1}=35M_{\rm{\odot}},m_{2}=30M_{\rm{\odot}}$, and $m_{1}=85M_{\rm{\odot}},m_{2}=66M_{\rm{\odot}}$, respectively.
Based on the selected parameter ranges described above, our calculation results are presented in FIG. 3. The purple, orange and green lines represent O3 PSD, Aplus PSD, and ET PSD, respectively. The solid, dashed and dot-dashed lines correspond to three mass combinations, i.e., $m_{1}=m_{2}=20M_{\rm{\odot}}$, $m_{1}=35M_{\rm{\odot}},m_{2}=30M_{\rm{\odot}}$, and $m_{1}=85M_{\rm{\odot}},m_{2}=66M_{\rm{\odot}}$, respectively. For the three PSD scenarios considered, we have computed the variation of the SNR with luminosity distance for three distinct mass combinations. Subsequently, based on the results obtained in this subsection, we will proceed to calculate the distribution of the mismatch.
III.2 Mismatch Arising from AGN Lensing
In this subsection, We first introduce the distribution of the mismatch on the AGN disk for the O3 PSD case. Then in the same way, we apply the calculations for the Aplus PSD and ET PSD cases.
III.2.1 Mismatch distribution: with O3 PSD
A schematic illustration of the lensing system is shown in FIG. 1 ($a$). As indicated by Eq. (15), the mismatch is determined by parameters $a$, $\phi$, $\iota$, and $M_{\rm{Lz}}$. The lensing related parameter ranges adopted in our analysis are as follows: $a\in[500R_{\rm S},5000R_{\rm S}]$, $\phi\in[-25^{\circ},\,25^{\circ}]$ and $\iota\in[65^{\circ},\,115^{\circ}]$. This parameter configuration is based on the following considerations: BHs embedded in the disk migrate inward and form binaries at the location of the migration trap [107], eventually merging through various mechanisms within this region [98, 99]. Reference [115] simulated this process, and their results indicate that the distribution of BBH pair-up locations depends on the AGN disk profile and properties, as well as the mass of the central SMBH [24]. In the majority of cases, BBHs form at the migration trap, which is typically located at distances ranging from several hundred to several thousand Schwarzschild radii. Furthermore, since the distribution of BHs in the accretion disk primarily follows a radial power-law profile [120, 69], we treat the inclination angle $\iota$ and the orbital phase $\phi$ as free parameters in our calculation.
We employ the PyCBC python package [16, 15, 36, 114, 85, 84] to compute the mismatch between the lensed waveform and the corresponding unlensed template waveform [86, 114]. The unlensed GW waveform template are generated using the IMRPhenomXPHM waveform model [94]. In the calculation of the mismatch, except for the change in the source mass, all other parameters are set to their default values. The mismatch distribution is shown in FIG. 4.